OpenStudy (anonymous):

Write a system of two equations using x and y to represent the following problem. A chemist is mixing two solutions together. The chemist wants 100 gallons when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use?

OpenStudy (anonymous):

Substitute the above into equation 2: .24x + .50y = .36(100) .24x + .50(100-x) = .36(100) .24x + 50-.50x = 36 50-.26x = 36 -.26x = -14 x = 53.85 gallons (of 24% solution) . to find 50% solution substitute above into equation 1: x+y = 100 53.85+y = 100 y = 46.15 gallons

OpenStudy (anonymous):

x=700/13 gallons

OpenStudy (mertsj):

x+y=100 .24x+.50y=.36(100)